OpenStudy (anonymous):

T=(N-3.99) / 0.17 n is the number of women enrolled ( in millions) t represent the number of years Use the new formula in part b to estimate the year in which women's enrollment will reach 13 million

OpenStudy (anonymous):

I did that but it came out weird numbers instead of the year

OpenStudy (astrophysics):

What does the 3.99 mean, what does 0.17 mean

OpenStudy (astrophysics):

What is part a

OpenStudy (astrophysics):

You need to give us all the information

OpenStudy (anonymous):

well the original formula is n=0.17t + 3.99

OpenStudy (anonymous):

but they asked me to solve for t using the equation n=0.17t + 3.99 in part b

OpenStudy (astrophysics):

This is all vague, you're just throwing out formulas and not explaining what they are

OpenStudy (anonymous):

okay let me send a pic

OpenStudy (astrophysics):

Alright

OpenStudy (astrophysics):

Oh I see...

OpenStudy (astrophysics):

n = 13

OpenStudy (astrophysics):

Because it states its in millions

OpenStudy (astrophysics):

And initially your year is 1970, I can't really read it as it's a very blurry pic

OpenStudy (astrophysics):

Yeah so all you have to do is plug in n = 12 as it's asking for that, and then you will get a number of years, and you add that to your initial year which is 1970 that will give you the year where 12 million woman were in college.

OpenStudy (astrophysics):

You can test with the other results as well to make sure, try n = 10, you will see you get 2005.

OpenStudy (astrophysics):

Make sense?

OpenStudy (anonymous):

oh okay I got it. thank you very much

OpenStudy (astrophysics):

For example \[N=0.17t+3.99\] lets make n = 10 this means 10 million woman \[10=0.17t+3.99 \implies t = \frac{ 10-3.99 }{ 0.17 } = 35\] 1970+35 = 2005

OpenStudy (anonymous):

yea I got 2017 for the answer. thank you again ~

OpenStudy (astrophysics):

Np

OpenStudy (astrophysics):

That's for 12 years, ok I see they are asking for 13 years in part C so do that as well haha

OpenStudy (anonymous):

yea I noticed lol